# Elias’ Encoding from Lagrangians and Renormalization

### v5

@@ Abstract
In the present paper we give a derivation of Elias’ Omega code from physics principles by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions.
Starting from a Lagrangian that minimizes average codelength under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated log-sum length, asymptotically up to an additive constant.
This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and entropy relations familiar from statistical physics.
@@ 2.1 Lagrangian Optimization for Codelength
Here, if $x\notin\mathcal{C}$ is not a codeword, we assume $\ell(x)=+\infty$. Let us note that for countable codes $p(x)$ is a linear combination of delta-functions, and the integrals above become the usual discrete sums. More details on prefix-free codes with mixed source distributions are provided in Appendix 3.
@@ 2.2 Renormalizing the Codeword Distribution
is the Shannon entropy [16] of $\mathcal{C}$. On the other hand, Shannon’s classical lower bound $\mathbb{E}[\ell]\geq H(p)$ also holds (Appendix 3.5).
@@ 2.4 Passing to the Discrete Encoding
Acknowledgments
3 Appendix: Continuous codes and Quantization
@@ 3.4 Heisenberg-style and Boltzmann-style relations
3.5 Appendix: Shannon’s Entropy Lower Bound
$p(dx)=\sum_{w\in W}\delta_{w}(dx)+\rho(x)dx,$
where $W\cap\mathrm{supp}\,\rho=\emptyset$.
A length function $\ell:\mathcal{C}\to[0,\infty]$ is admissible if
$\int e^{-\ell(x)}\,\nu(dx)\;\leq\;1.$
Then the entropy is
We assume that $\mathcal{C}$ satisfies the Kraft condition:
$K=\int e^{-\ell(x)}\,\nu(dx)\leq 1,$
and call
$q(x)=\frac{e^{-\ell(x)}}{K}$
the Kraft distribution.
The latter is equivalent to
$\ell(x)=-\ln q(x)-\ln K,$
and thus we obtain
$=H(p)+D_{\mathrm{KL}}(p\|q)-\ln Z\;\geq\;H(p),$
Thus, we have an analogue of Shannon’s lower entropy bound
$\mathbb{E}_{p}[\ell]\ \geq\ H(p),$
with equality if and only if $K=1$ and $e^{-\ell}=\frac{dp}{d\nu}$ ($\nu$-a.e.).

### v6

# and Quantization in Prefix Coding

@@ Abstract
We develop a statistical mechanics framework for prefix coding based on variational principles, renormalization, and quantization. A Lagrangian formulation of the Minimal Entropy principle under the Kraft–McMillan constraint yields a Gibbs-type implied distribution and completeness of the optimal code. A renormalization operator acting on codeword distribution laws produces a coarse-graining flow whose fixed points have iterated-log structure; discrete quantizations of these fixed points include Elias’ $\omega$ code as a special case.
Extending the theory to mixed discrete–continuous source laws, we show how continuous codelength functions can be quantized into countable prefix codes and derive resolution-adjusted entropy bounds together with Heisenberg-type and Boltzmann-type relations. This provides a unified and physically motivated view of universal coding, with Elias’ $\omega$ code as a guiding example.
@@ 2.1 Lagrangian Optimization for Codelength
Here, if $x\notin\mathcal{C}$ is not a codeword, we assume $\ell(x)=+\infty$. Let us note that for countable codes $p(x)$ is a linear combination of delta-functions, and the integrals above become the usual discrete sums. More details on prefix-free codes with mixed source distributions are provided in Section 4.
@@ 2.2 Renormalizing the Codeword Distribution
is the Shannon entropy [16] of $\mathcal{C}$. On the other hand, Shannon’s classical lower bound $\mathbb{E}[\ell]\geq H(p)$ also holds (cf. Section 4.4).
@@ 2.4 Passing to the Discrete Encoding
3 Discussion
The case of Elias’ $\omega$ encoding shows that it can be fruitful to look at an intrinsically discrete object through a continuous lens. Elias’ $\omega$ then appears as a particular discretization or, rather, quantization of a continuous distribution obtained as a fixed point of a flow, by replacing $x$ with the binary length $\beta(n)$ and terminating the flow when the argument drops to $1$.
In the derivation above we never used the fact that codewords are integers until the very last step: the renormalization flow and the fixed-point equation were all formulated for a continuous codelength function on $1,\infty)$ endowed with a probability measure $p(dx)$.
From this point of view, the continuous formulation is not a cosmetic reformulation of a known code, but a device that explains why such a code exists and why it has a very specific mix of properties. The Lagrangian construction identifies an implied Gibbs measure and forces completeness by saturating the Kraft–-McMillan inequality; the renormalization operator selects a fixed point whose codelength grows like an iterated sum of logarithms. The discrete Elias $\omega$ code is then recovered as a quantized version of this continuous picture.
This perspective suggests a more general discussion continued below by introducing continuous codes and their quantization, and by deriving the corresponding Kraft inequalities, entropy bounds, and resolution–complexity tradeoffs.
4 Continuous codes and Quantization
@@ 4.3 Prefix-free Quantized Codes
4.4 Shannon’s Entropy Lower Bound
$p(dx)=\sum_{w\in W}\delta_{w}(dx)+\rho(x)dx,$
where $W\cap\mathrm{supp}\,\rho=\emptyset$.
A length function $\ell:\mathcal{C}\to[0,\infty]$ is admissible if
$\int e^{-\ell(x)}\,\nu(dx)\;\leq\;1.$
Then the entropy is
We assume that $\mathcal{C}$ satisfies the Kraft condition:
$K=\int e^{-\ell(x)}\,\nu(dx)\leq 1,$
and call
$q(x)=\frac{e^{-\ell(x)}}{K}$
the Kraft distribution.
The latter is equivalent to
$\ell(x)=-\ln q(x)-\ln K,$
and thus we obtain
$=H(p)+D_{\mathrm{KL}}(p\|q)-\ln Z\;\geq\;H(p),$
Thus, we have an analogue of Shannon’s lower entropy bound
$\mathbb{E}_{p}[\ell]\ \geq\ H(p),$
with equality if and only if $K=1$ and $e^{-\ell}=\frac{dp}{d\nu}$ ($\nu$-a.e.).
@@ 4.5 Heisenberg-style and Boltzmann-style relations
5 Conclusion
Using Elias’ $\omega$ encoding as a guiding example, we have developed a general framework that links prefix coding with ideas and tools from statistical mechanics.
On the coding side, we start from a variational formulation of optimal codelengths under the Kraft–McMillan constraint. Then we show that the optimal solution automatically yields a complete code whose implied probability law coincides with the source distribution.
On the physics side, this variational problem naturally produces a Gibbs-style measure on codewords and a partition-function-like Kraft sum.
We then introduce a renormalization operator acting on codeword distributions, and the associated flow on codelengths. Its fixed point exhibits the characteristic iterated-log growth familiar from Elias’ $\omega$ code.
The flow has a relatively wide basin of attraction and preserves asymptotic optimality up to an additive constant, showing that the scale-invariant structure behind Elias’ construction is not an accidental but rather a robust feature of a larger family.
Finally, we extend the Kraft–McMillan framework to mixed discrete–continuous source laws and make the passage back to practical discrete codes explicit via quantization. In this setting, the average codelength decomposes into a resolution-adjusted entropy term plus non-negative corrections capturing within-cell structure, code-tree mismatch, and Kraft’s slack.
The above decomposition leads to an uncertainty-type relation that trades off resolution against coding complexity, and to a Boltzmann-style identity in which the optimal average codelength measures the logarithm of an effective number of encodable macrostates.
Acknowledgments